OpenStudy (anonymous):

Write an equation of a parabola with its vertex at the origin and its directrix at y = –2.

OpenStudy (dumbcow):

\[y =a(x-h)^{2} +k\] equation for directrix \[y = k -\frac{1}{4a}\] h=0 k=0 solve for a

OpenStudy (anonymous):

is it -2y^2

OpenStudy (anonymous):

???

OpenStudy (dumbcow):

plug in -2 for y \[-2 = -\frac{1}{4a}\]

OpenStudy (anonymous):

A.) y=1/8y^2 B.) x=-2y^2 C.) x=1/8x^2 D.) y=8x^2

OpenStudy (anonymous):

A ??

OpenStudy (anonymous):

thats what i got is it right?

OpenStudy (anonymous):

this is the y=x^2 type because your directrix is horizontal. the distance from vertex to directrix is 2 so a=2. since your vertex is at the origin, your equation should be y=(1/8)x^2

OpenStudy (dumbcow):

yes